OpenStudy (anonymous):

3lnx^2 + 4ln3 write as a single logarithm

OpenStudy (mertsj):

\[alog (n)=\log(n)^a\]

OpenStudy (mertsj):

Use that property on both terms.

OpenStudy (mertsj):

\[\log(m)+\log(n)=\log(mn)\]

OpenStudy (mertsj):

Then use that property on the two terms to get a single term.

OpenStudy (mertsj):

@precal You are incorrect.

OpenStudy (mertsj):

\[3\ln x^2=\ln x^6\]

OpenStudy (precal):